Consider the following data to validate the hypothesis that there is significant difference


between two populations.


Sample size


Mean


Standard deviation


20


25


5


20


20


6



1
Expert's answer
2021-07-01T11:00:27-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu_1=\mu_2"

"H_1:\mu_1\not=\mu_2"

This corresponds to a two-tailed test, and a z-test for two means, with known population standard deviations will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a two-tailed test is "z_c=1.96." The rejection region for this two-tailed test is 

"R=\{z:|z|>1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x_1}-\bar{x_2}}{\sqrt{\sigma_1^2/n_1+\sigma_2^2/n_2}}"

"=\dfrac{25-20}{\sqrt{5^2/20+6^2/20}}=2.863"

Since it is observed that "|z|=2.863>1.96=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value is "p=2P(Z>2.863)=0.0042," and since "p=0.0042<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu_1" is different than "\mu_2," at the "\alpha=0.05" significance level.